OpenStudy (emilyjones284):

ill give a medal (:

OpenStudy (emilyjones284):

\[\sqrt{x+2}=10-x\]

OpenStudy (anonymous):

Start by squaring both sides to remove the radical: \[x+2=100 -20x +x^2\] Subtract x and 2 from both sides: \[0=x^2 -21x +98\] From here, you could try to factor, but I would just plug into the quadratic formula. If you'd like more assistance with this let me know.

OpenStudy (emilyjones284):

do you know if the answer would be 7 or 14 or both because im not really sure how to check for extraneous solutions

OpenStudy (anonymous):

Plug it back in the original equation, if it gives an untrue statement=extraneous

OpenStudy (anonymous):

For extraneous solutions just substitute back the values in the original equation. If you get any absurd thing, its a extr. solution

OpenStudy (emilyjones284):

okay thanks